OpenStudy (anikate):

Riddle: What is 1/3 * 3?

OpenStudy (anikate):

Get a medal if u answer

OpenStudy (aaaa):

1

OpenStudy (anikate):

.wrong wanna know why?

OpenStudy (anikate):

u got close though

OpenStudy (anonymous):

1/9

OpenStudy (anikate):

nope

OpenStudy (anonymous):

0

OpenStudy (anikate):

nope

OpenStudy (anonymous):

infinity

OpenStudy (anikate):

nope

OpenStudy (karatechopper):

1/27

OpenStudy (anikate):

nope

OpenStudy (anikate):

say pass if u all give up

OpenStudy (aaaa):

3/3

OpenStudy (anikate):

very close but no

OpenStudy (karatechopper):

http://www.google.com/search?aq=f&ix=icb&sourceid=chrome&ie=UTF-8&q=What+is+1%2F3+*+3%3F <--- google must be wrong then

OpenStudy (aaaa):

0.99999999999999999999999999999999999999999999999999999

OpenStudy (karatechopper):

4/3

OpenStudy (anikate):

no google is correct in its own way

OpenStudy (karatechopper):

ok

OpenStudy (anikate):

I'm asking a riddle

OpenStudy (aaaa):

0.99999999999999999999999999999999999999999999999999999

OpenStudy (anikate):

wanna know the answer

OpenStudy (anikate):

correct aaa

OpenStudy (karatechopper):

looks like a question and yes

OpenStudy (karatechopper):

how?

OpenStudy (aaaa):

swagg

OpenStudy (anonymous):

the answer is 1 .. in math terms anyways

OpenStudy (anikate):

lolz

OpenStudy (anikate):

correct kmila

OpenStudy (aaaa):

im smarter than yalls hahahahahahahahahahahahahahhahahahaha

OpenStudy (anikate):

1/3=.3333333 so, .3333333*3=.99999999999999999999999999999999999=9/10

OpenStudy (karatechopper):

o

OpenStudy (anikate):

yup. :D

OpenStudy (anonymous):

Don't that many 9's if you times that many 3s

OpenStudy (anikate):

i no

OpenStudy (asnaseer):

1/3 * 3 = 1 if you do the multiplication symbolically (1/3) * 3 = 0.999... otherwise BTW: 3*0.333... = 0.999... does not equal 9/10. 9/10 = 0.9